### v2

# Endpoint variation and jump inequalities for rough singular integrals

Abstract.
In this article, we prove weak type $(1,1)$ bounds for the variation and jump operators corresponding to the family of truncations of singular integrals with rough kernels. This resolves an open question raised by Jones, Seeger and Wright (Trans. Amer. Math. Soc. (2008)). Moreover, as an immediate consequence of the variational estimate, we recover the weak type $(1,1)$ boundedness of the maximal truncation operator corresponding to singular integrals with rough kernels.
Key words and phrases:
Variation inequalities, jump inequalities, weak type estimate, maximal operator
2020 Mathematics Subject Classification:
Primary 42B25, 42B20
1. Introduction
The study of variational and jump inequalities is a recurrent theme in harmonic analysis and ergodic theory. A vast amount of research, see for example [Lep1976, Bourgain1989, JKRW1998, JRW2001, JRW2003, CJRRW2000, CJRW2003, JSW2008] have been devoted to the study of such inequalities. This article is concerned with the endpoint estimates for variational and jump operators for the family of truncations of singular integral operators with rough kernels defined as follows:
$T_{\Omega,\epsilon}f(x)=\int_{|x-y|>\epsilon}\Omega\left(\frac{x-y}{|x-y|}\right)|x-y|^{-d}f(y)dy,$
where $\Omega\in L\log L({\mathbb{S}}^{d-1})$ with $\int\Omega(\theta)d\sigma(\theta)=0$.
Calderón and Zygmund [CZ1956] showed that the operators $T_{\Omega}$ and its maximal variant $T_{\Omega}^{*}$ defined as
$T_{\Omega}f(x)=\lim_{\epsilon\to 0}T_{\Omega,\epsilon}f(x),\quad T_{\Omega}^{*}f(x)=\sup_{\epsilon>0}|T_{\Omega,\epsilon}f(x)|,$
are bounded in $L^{p}({\mathbb{R}}^{d})$ for $1<p<\infty$. We also refer to [DR1986] for an alternate proof involving a double dyadic analysis and Fourier transform estimates. At the endpoint $p=1$, Christ and Rubio de Francia [CR1988] and independently Hofmann [Hofmann1988] showed that $T_{\Omega}$ maps $L^{1}({\mathbb{R}}^{d})$ to $L^{1,\infty}({\mathbb{R}}^{d})$ for $d=2$. The weak $(1,1)$ boundedness for $T_{\Omega}$ was extended for all dimensions $d\geq 2$ by Seeger [Seeger1996] using a microlocal analysis of the kernel. The boundedness of the maximal operator $T_{\Omega}^{*}$ at the endpoint $p=1$ remained open until very recently. Honzík [Honzik2020] showed that $T_{\Omega}^{*}$ maps the space $L(\log\log L)^{2+\epsilon}({\mathbb{R}}^{d})$ to $L^{1,\infty}({\mathbb{R}}^{d})$ locally for $\epsilon>0$ and $\Omega\in L^{\infty}({\mathbb{S}}^{d-1})$. A further improvement was given by the first author and Mohanty in [BM2023], where they showed that $T_{\Omega}^{*}$ maps the larger space $L\log\log L({\mathbb{R}}^{d})$ to $L^{1,\infty}({\mathbb{R}}^{d})$ for $\Omega\in L\log L({\mathbb{S}}^{d-1})$. Recently, Lai [Lai2025] showed that $T_{\Omega}^{*}$ indeed maps $L^{1}({\mathbb{R}}^{d})$ to $L^{1,\infty}({\mathbb{R}}^{d})$.
In this paper, we are concerned with the variation and jump inequalities corresponding to the families of truncations of singular integrals with rough kernels. For $1<q\leq\infty$, consider the $q$-variations associated to a family of functions $\{F_{t}:t\in\mathcal{I}\}$ defined as follows
$V_{q}(F_{t}(x):t\in\mathcal{I})=\begin{cases}\sup\limits_{N\in{\mathbb{N}}}\sup\limits_{\begin{subarray}{c}t_{1}<t_{2}<\cdots t_{N}\\ t_{i}\in\mathcal{I}\end{subarray}}\left(\sum_{i=1}^{N-1}|F_{t_{i+1}}(x)-F_{t_{i}}(x)|^{q}\right)^{\frac{1}{q}},&1<q<\infty\\ \sup\limits_{t_{1},t_{2}\in\mathcal{I}:t_{1}<t_{2}}|F_{t_{1}}(x)-F_{t_{2}}(x)|,&q=\infty.\end{cases}$
The $\lambda$-jump corresponding to the family of functions $\{F_{t}:t\in\mathcal{I}\}$ is defined as follows
[…]
$\text{such that}|F_{t_{i+1}}(x)-F_{t_{i}}(x)|>\lambda\Big\}.$
We refer the reader to [JSW2008] for more details about the notion of variational and jump inequalities. For notational convenience, we write $V_{q}(F_{t}(x))=V_{q}(F_{t}(x):t\in(0,\infty))$ for the special case $\mathcal{I}=(0,\infty)$. Similarly, we set $N_{\lambda}(F_{t}(x))=N_{\lambda}(F_{t}(x):t\in(0,\infty))$.
Campbell, Jones, Reinhold and Wierdl [CJRRW2000, CJRW2003] investigated the $L^{p}$-bounds of variational and jump inequalities for the family $\{T_{\Omega,\epsilon}:\;\epsilon\in(0,\infty)\}$. In particular, they proved the following bounds for $\Omega\in L\log L({\mathbb{S}}^{d-1})$ and $1<p<\infty$:
$\left\|V_{q}(T_{\Omega,\epsilon}f)\right\|_{p}\lesssim\|f\|_{p},\quad\text{for}\;q>2,$ (1.1)
$\left\|\lambda(N_{\lambda}(T_{\Omega,\epsilon}f))^{\frac{1}{\rho}}\right\|_{p}\lesssim\|f\|_{p},\quad\text{for}\;\rho>2.$ (1.2)
Jones, Seeger and Wright [JSW2008] established the strong type inequality (1.2) at the critical case $\rho=2$ for $\Omega\in L^{r}({\mathbb{S}}^{d-1}),\;1<r\leq\infty$. The estimate (1.2) for $\rho=2$ and $\Omega\in L\log L({\mathbb{S}}^{d-1})$ was obtained by Ding, Hong and Liu [DHL2017].
In [CJRRW2000, CJRW2003], the following weak type $(1,1)$ bounds, when $\Omega$ is a Lipschitz function, were also established.
$\left|\left\{x\in{\mathbb{R}}^{d}:\;V_{q}(T_{\Omega,\epsilon}f(x))>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha},\quad\text{for}\;q>2,$ (1.3)
$\left|\left\{x\in{\mathbb{R}}^{d}:\;\lambda(N_{\lambda}(T_{\Omega,\epsilon}f(x)))^{\frac{1}{\rho}}>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha},\quad\text{for}\;\rho>2.$ (1.4)
In [JSW2008] the weak type $(1,1)$ inequality (1.4) was proved at the critical case $\rho=2$ for $\Omega$ being a Lipschitz function. In the same paper, the question of weak type $(1,1)$ boundedness of the underlying operator in the absence of any smoothness assumption on $\Omega$ was posed and left open. In this paper, we resolve this question affirmatively; this constitutes the main result of our work.
Theorem 1.1.
Let $\Omega\in L\log L({\mathbb{S}}^{d-1})$. Then, for all $\alpha>0$, we have
$\left|\left\{x\in{\mathbb{R}}^{d}:\lambda[N_{\lambda}(T_{\Omega,\epsilon}f(x))]^{\frac{1}{2}}>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha}.$ (1.5)
Moreover for $q>2$, we have
$\left|\left\{x\in{\mathbb{R}}^{d}:\;V_{q}(T_{\Omega,\epsilon}f(x))>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha}.$ (1.6)
By a standard limiting argument, Theorem 1.1 recovers the weak $(1,1)$ boundedness of $T_{\Omega}^{*}$ proved recently by Lai [Lai2025].
Methodology of the proof of Theorem 1.1:
We first note that by [MSZ2020, Lemma 2.3], the weak type $(1,1)$ estimate (1.6) for the $q$-variation operator will follow from the corresponding weak type estimate (1.5) involving the jump function. Further, it is well known that in order to prove (1.5), it is enough to consider the short and long jump operators separately via the inequality (2). In both the cases, the variational Rademacher-Menshov theorem is a key ingredient. This result was first obtained in [LL2012]. We will use the following version of the variational Rademacher-Menshov theorem from [DOP2017, Lemma 7.2].
Lemma 1.2 ([DOP2017]).
Let $\{F_{i}\}_{i=1}^{N}$ be a sequence of functions on a measure space $X$ such that for all $\varepsilon_{i}\in\{-1,0,1\}$, we have
$\left\|\sum_{i=1}^{N}\varepsilon_{i}F_{i}\right\|_{2}\leq B.$
Then, the following estimate holds
$\left\|V_{2}\left(\sum_{i=1}^{n}F_{i}:1\leq n\leq N\right)\right\|_{2}\lesssim(\log N)B.$
The second ingredient consists of microlocal estimates of Seeger [Seeger1996]. Our task is to combine the $L^{1}$ and $L^{2}$ estimates of Seeger [Seeger1996] into a single $L^{2}$ estimate. This step relies on a careful interpolation argument, motivated by the approach in [Honzik2020]. We provide the details of this estimate in Appendix A.
The main distinction between the short and long jumps is the presence of multiscale analysis in the estimate for the long jump. The estimate for the short jumps relies on a single scale argument reminiscent of the ideas developed in [CJRW2003]. In contrast, the estimates for the long jumps are based on an algorithm that controls the interaction of scales, pioneered in the work of Krause and Lacey [KL2018, KL2020] for obtaining sparse bounds for maximal truncations of oscillatory Calderón-Zygmund operators. We also refer to [IKM2025] for a simplified approach to proving endpoint estimates for such maximal oscillatory operators in dimension one.
In the process of adapting the ideas of [KL2020], we streamline certain aspects of their proof. First, we do away with the recursion argument employed there. Second, we introduce an effective decomposition of the bad functions based on the shifted dyadic grids. This allows us to use the $L^{2}$-bounds (Lemma A.2) as a blackbox instead of arduously reproving similar estimates for a localised version of the truncations of rough singular integrals. As a consequence, this approach provides us with a new and simpler proof of the weak $(1,1)$-boundedness of $T_{\Omega}^{*}$. The proof of Theorem 1.1 (1.5) is given in Section 2.
2. Proof of Theorem 1.1 (1.5)
Let $f\in L^{1}({\mathbb{R}}^{d})$ and $\alpha>0$. By the Calderón-Zygmund decomposition, applied to $f$ at height $\alpha$, there exists a family $\mathcal{Q}$ of disjoint dyadic cubes such that
  1. (1)
    $f=g+b$ with $b=\sum_{s\in{\mathbb{Z}}}B^{s}$, where $B^{s}=\sum_{Q\in\mathcal{Q}:|Q|=2^{sd}}b_{Q}$,
  2. (2)
    $\|g\|_{\infty}\lesssim\lambda$ and $\|g\|_{1}\leq\|f\|_{1}$,
  3. (3)
    $\int b_{Q}(x)dx=0$ and $\operatorname{supp}(b_{Q})\subseteq Q$,
  4. (4)
    $\|b_{Q}\|_{1}\lesssim\alpha|Q|$ and $\sum_{Q\in\mathcal{Q}}|Q|\lesssim\frac{1}{\alpha}$. In particular, this implies $\sum_{s}\|B^{s}\|_{1}\lesssim\|f\|_{1}$.
We define the exceptional set $E=\cup_{Q\in\mathcal{Q}}4Q$. Thus $|E|\lesssim\frac{\|f\|_{1}}{\alpha}$. By the $L^{2}$-boundedness [DHL2017, Theorem 1.2 (ii)] of the $\lambda$-jump operators associated to $\{T_{\Omega,\epsilon}:\epsilon>0\}$, we have
$|\{x\in{\mathbb{R}}^{n}:\lambda[N_{\lambda}(T_{\Omega,\epsilon}g(x):\epsilon\in(0,\infty))]^{\frac{1}{2}}>\alpha\}|\lesssim\frac{1}{\alpha^{2}}\|g\|_{2}^{2}\lesssim\frac{\|f\|_{1}}{\alpha}.$
Therefore it remains to estimate for the bad part $b$. Let $\psi$ be a Schwartz function supported in the annulus $\{\frac{1}{4}\leq|y|\leq 1\}$ and $\sum_{j\in{\mathbb{Z}}}\psi(2^{-j}y)=1$. We decompose the kernel as follows
$K(y)=\Omega\left(\frac{y}{|y|}\right)|y|^{-d}=\sum_{j\in{\mathbb{Z}}}\Omega\left(\frac{y}{|y|}\right)|y|^{-d}\psi(2^{-j}y)=\sum_{j\in{\mathbb{Z}}}K_{j}(y).$
We note that by the support of the kernels $K_{j}$, it follows that
$\sum_{s<5}\sum_{j\in{\mathbb{Z}}}(K_{j}\chi_{|\cdot|>\epsilon})*B^{j-s}(x)=0,\quad\text{for}\;x\in E^{c}.$
Thus, we have that
$\lambda[N_{\lambda}(T_{\Omega,\epsilon}b(x):\epsilon\in(0,\infty))]^{\frac{1}{2}}\leq\lambda\left[N_{\lambda}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(K_{j}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}.$
For each $s\geq 5$, we write $K_{j}=S_{j}^{s}+H_{j}^{s}$, where
$S_{j}^{s}(y)$ $=K_{j}(y)\chi_{\left\{\Omega\left(\frac{y}{|y|}\right)>2^{\delta_{0}s}\|\Omega\|_{1}\right\}}(y),$
$H_{j}^{s}(y)$ $=K_{j}(y)\chi_{\left\{\Omega\left(\frac{y}{|y|}\right)\leq 2^{\delta_{0}s}\|\Omega\|_{1}\right\}}(y),$
for some fixed $\delta_{0}>0$. Thus, we obtain
$\lambda[N_{\lambda}(T_{\Omega,\epsilon}b:\epsilon\in(0,\infty))]^{\frac{1}{2}}(x)\leq$ $\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$
$+\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon>0\right)\right]^{\frac{1}{2}}.$
For the term involving operators $S_{j}^{s}$ in the above, by the definition of $\lambda$-jumps and variation, we have that
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$
$\lesssim$ $V_{1}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
$\leq$ $|S_{j}^{s}|*|B^{j-s}|(x).$
Hence, we use Chebyshev’s inequality to obtain
$\left|\left\{x\in E^{c}:\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}>\frac{\alpha}{2}\right\}\right|$
$\lesssim$ $\frac{1}{\alpha}\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}\|S_{j}^{s}\|_{1}\|B^{j-s}\|_{1}$
$\lesssim$ $\frac{1}{\alpha}\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}\|B^{j-s}\|_{1}\int_{\Omega(\theta)>2^{\delta_{0}s}\|\Omega\|_{1}}\Omega(\theta)d\theta$
$\lesssim$ $\frac{1}{\alpha}\|\Omega\|_{L\log L}\|f\|_{1}.$
For the term involving operators $H_{j}^{s}$, we dominate the $\lambda$-jumps by the sum of short and long $\lambda$-jumps. This is achieved via a standard argument; for a proof, see [JSW2008, Lemma 1.3]. As a consequence, we obtain the following pointwise inequality:
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$
$\lesssim$ $~\mathcal{S}_{2}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
$+\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>2^{k}})*B^{j-s}(x):k\in{\mathbb{Z}}\right)\right]^{\frac{1}{2}},$ (2.1)
where $\mathcal{S}_{2}$ is the short jump operator defined as
$\mathcal{S}_{2}(F_{\epsilon}:\epsilon>0)(x)=\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}(F_{\epsilon}:\epsilon\in[2^{k-1},2^{k}])(x)|^{2}\right)^{\frac{1}{2}}.$
The second term in the expression (2) is known as the long jump function.
2.1. Estimate for short jumps
Since the support of $H_{j}^{s}$ is contained in the annulus $\{2^{j-2}\leq|y|\leq 2^{j}\}$, the estimate of $\mathcal{S}_{2}(T_{\Omega,\epsilon}b:\epsilon\in(0,\infty))$ is reduced to proving
$\left|\left\{x\in E^{c}:\sum_{s\geq 5}\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}((H^{s}_{k+j}\chi_{|\cdot|>\epsilon})*B^{k+j-s}(x):\epsilon\in[2^{k-1},2^{k}])|^{2}\right)^{\frac{1}{2}}>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha},$
for $j=0,1$. We will prove the inequality for $j=0$, the other estimate is similar. By Chebyshev’s and triangle inequality, it is enough to prove the following for $k\in{\mathbb{Z}}$ and $s\geq 5$,
$\left\|V_{2}((H_{k}^{s}\chi_{|\cdot|>\epsilon})*B^{k-s}:\epsilon\in[2^{k-1},2^{k}])\right\|_{2}^{2}\lesssim 2^{-\delta^{\prime}s}\alpha\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)n}}\|b_{Q}\|_{1},$
for some $\delta^{\prime}>0$. Recalling the definition of $V_{2}$, the above inequality is equivalent to
$\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}(x)|^{2}\right)^{\frac{1}{2}}\right\|^{2}\lesssim 2^{-\delta^{\prime}s}\alpha\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)d}}\|b_{Q}\|_{1},$
where $H^{s}_{k,E_{i}^{k}(x)}=H^{s}_{k}\chi_{\epsilon_{i-1}^{k}(x)\leq|\cdot|\leq\epsilon_{i}^{k}(x)}$ for a measurable choice of functions $N^{k}(x)$ and $\epsilon_{i}^{k}(x),\;i=0,\cdots,N^{k}(x)$ with $2^{k-1}\leq\epsilon_{0}^{k}(x)<\epsilon_{1}^{k}(x)<\cdots<\epsilon_{N}^{k}(x)(x)\leq 2^{k}$. We define the functions $B^{k-s}_{i,in}$ and $B^{k-s}_{i,bd}$ as follows
$B^{k-s}_{i,in}=\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:\\ |Q|=2^{(k-s)d}\\ (x-Q)\subseteq E_{i}^{k}(x)\end{subarray}}b_{Q},\qquad B^{k-s}_{i,bd}=\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:\\ |Q|=2^{(k-s)d}\\ (x-Q)\cap(E_{i}^{k}(x))^{c}\neq\phi\end{subarray}}b_{Q},$
for $E_{i}^{k}(x)=[\epsilon_{i-1}^{k}(x),\epsilon_{i}^{k}(x)]$. Thus, we have
$\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{-s}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$
$\lesssim$ $\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,in}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}+\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,bd}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$
$=:$ $I_{in}+I_{bd}.$
2.1.1. Estimate for $I_{bd}$:
We note that $|H^{s}_{k,E_{i}^{k}(x)}|\lesssim\|\Omega\|_{1}2^{\delta_{0}s}2^{-kd}$ and the maximum number of cubes $Q\in\mathcal{Q}$ with side length $2^{k-s}$ such that $(x-Q)\cap\partial E_{i}^{k}(x)\neq\phi$ is proportional to $2^{(s)(d-1)}$. Thus, we have
$|H^{s}_{k,E_{i}^{k}(x)}*B_{i,bd{{}^{-s}}}(x)|$ $\lesssim\|\Omega\|_{1}2^{\delta_{0}s}2^{-kd}\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:\\ |Q|=2^{(k-s)d}\\ (x-Q)\cap(E_{i}^{k}(x))^{c}\neq\phi\end{subarray}}\|b_{Q}\|_{1}$
$\lesssim\alpha\|\Omega\|_{1}2^{\delta_{0}s}\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:\\ |Q|=2^{(k-s)d}\\ (x-Q)\cap(E_{i}^{k}(x))^{c}\neq\phi\end{subarray}}|Q|$
$\lesssim\alpha\|\Omega\|_{1}2^{-(1-\delta_{0})s}.$
Therefore, it follows that
$I_{bd}$ $\lesssim\int\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,bd}(x)|^{2}dx$
$\lesssim\alpha\|\Omega\|_{1}2^{-(1-\delta_{0})s}\int\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,bd}(x)|dx$
$\lesssim\alpha\|\Omega\|_{1}2^{-(1-\delta_{0})s}\int\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}|*\left(\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)d}}|b_{Q}|\right)(x)dx$
$\lesssim\alpha\|\Omega\|_{1}^{2}2^{-(1-2\delta_{0})s}\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)d}}\|b_{Q}\|_{1}.$
2.1.2. Estimate for $I_{in}$:
We begin by noting that
$H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,in}(x)$ $=\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:\\ |Q|=2^{(k-s)d}\\ (x-Q)\subseteq E_{i}^{k}(x)\end{subarray}}K_{k}*b_{Q}(x)$
$=\sum_{m\in{\mathbb{Z}}^{d}}K_{k}*B^{k-s,m}_{i,in}(x),\;\text{where}$
$B^{k-s,m}_{i,in}=\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:|Q|=2^{(k-s)d},\\ Q\subseteq 2^{k}(m+[0,1]^{d})\\ (x-Q)\subset E_{i}^{k}(x)\end{subarray}}b_{Q},\quad\text{and}\quad B^{k-s,m}=\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:|Q|=2^{(k-s)d},\\ Q\subseteq 2^{k}(m+[0,1]^{d})\end{subarray}}b_{Q}.$
We note the above quantities are defined only for those admissible $m\in{\mathbb{Z}}^{d}$ for which atleast one such $Q$ exists. The following single scale estimate with decay in the parameter $s$ holds via Lemma A.2.
$\|K_{k}*\widetilde{B}^{k-s,m}\|_{2}^{2}\lesssim 2^{-(\delta-\delta_{0})s}\alpha\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:|Q|=2^{(k-s)d}\\ Q\subseteq 2^{k}(m+[0,1]^{d})\end{subarray}}\|b_{Q}\|_{1},$ (2.2)
where $\widetilde{B}^{k-s,m}=\sum\limits_{\begin{subarray}{c}Q\in\mathcal{Q}:|Q|=2^{(k-s)d}\\ Q\subseteq 2^{k}(m+[0,1]^{d})\end{subarray}}\epsilon_{Q}b_{Q}$ for any fixed choice of $\epsilon_{Q}\in\{-1,0,1\}$.
This allows us to use the variational Rademacher-Menshov Lemma 1.2. To do so, we first observe that by spatial localization of the bad cubes $Q$, the number of $b_{Q}$’s appearing in the definition of $B^{k-s,m}_{i,in}$ is atmost $2^{sd}$ for each fixed $m\in{\mathbb{Z}}^{d}$, $x\in{\mathbb{R}}^{d}$, and $i=1,\cdots,N^{k}(x)$. We also note that the sequence of functions $\{H^{s}_{k}*B^{k-s,m}_{i,in}(x)\}_{m\in{\mathbb{Z}}}$ have bounded overlap. Hence, we have
$I_{in}$ $\lesssim\int\sum_{i=1}^{N^{k}(x)}\left|\sum_{m\in{\mathbb{Z}}^{d}}H^{s}_{k}*B^{k-s,m}_{i,in}(x)\right|^{2}dx$
$\lesssim\sum_{m\in{\mathbb{Z}}^{d}}\int\sum_{i=1}^{N^{k}(x)}\left|H^{s}_{k}*B^{k-s,m}_{i,in}(x)\right|^{2}dx$
$\lesssim 2^{-(\delta-\delta_{0})s}s^{2}\alpha\sum_{m\in{\mathbb{Z}}^{d}}\sum_{\begin{subarray}{c}Q\in\mathcal{Q}:|Q|=2^{(k-s)d}\\ Q\subseteq 2^{k}(m+[0,1]^{d})\end{subarray}}\|b_{Q}\|_{1}$
$\lesssim 2^{-(\delta-\delta_{0})s}s^{2}\alpha\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)d}}\|b_{Q}\|_{1},$
where we used Lemma 1.2 in the second last step. Choosing $\delta^{\prime}<\min\{1-2\delta_{0},(\delta-\delta_{0})/2\}$ concludes the proof for short jumps.
2.2. Estimate for long jumps
In this section we prove the weak $(1,1)$ bounds for the long jumps. We note that
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>2^{k}})*B^{j-s}:k\in{\mathbb{Z}}\right)\right]^{\frac{1}{2}}(x)$
$\lesssim$ $\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}$
$+\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}(H_{k}^{s}\chi_{2^{k-1}\leq|.|\leq 2^{k}})*B^{k-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}$
$=:$ $L_{1}+L_{2}$
2.2.1. Estimate for $L_{2}$:
The estimate for $L_{2}$ is comparatively easier as there is no interaction between scales $k$. Indeed, by Chebyshev’s and triangle inequality, we have
$\left|\left\{\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}(H_{k}^{s}\chi_{2^{k-1}\leq|.|\leq 2^{k}})*B^{k-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}>\frac{\alpha}{4}\right\}\right|$
$\lesssim$ $\frac{1}{\alpha^{2}}\left(\sum_{s\geq 5}\left(\sum_{k\in{\mathbb{Z}}}\|(H_{k}^{s}\chi_{2^{k-1}\leq|.|\leq 2^{k}})*B^{k-s}\|_{2}^{2}\right)^{\frac{1}{2}}\right)^{2}$
$\lesssim$ $\frac{\|\Omega\|_{1}}{\alpha}\left(\sum_{s\geq 5}2^{-(\epsilon-\delta_{0})s/2}\left(\sum_{k\in{\mathbb{Z}}}\|B^{k-s}\|_{1}\right)^{\frac{1}{2}}\right)^{2}$
$\lesssim$ $\frac{1}{\alpha}\|\Omega\|_{1}\|f\|_{1},$
where we have used Lemma A.2 in the second last step.
2.2.2. Estimate for $L_{1}$:
For the estimate of $L_{1}$, we note that
$\left|\left\{\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}>\frac{\alpha}{4}\right\}\right|$
$\leq\left|\left\{\sum_{s\geq 5}V_{2}\left(\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\right)(x)>\frac{\alpha}{4}\right\}\right|$
$\leq\sum_{s\geq 5}\left|\left\{V_{2}\left(\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\right)(x)>{c_{1}2^{-\delta_{1}s}\alpha}\right\}\right|,$
where $c_{1},\delta_{1}$ are fixed constants such that $c_{1}\sum_{s\geq 5}2^{-\delta_{1}s}=\frac{1}{4}$.
We will work with a fixed $s\geq 5$. We set $K_{j,m}=2^{j}([0,1)^{d}+m)$. As earlier, we write
$H_{j}^{s}*B^{j-s}=\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m},\quad\text{where}\;B^{j-s,m}=\sum_{|Q|=2^{(j-s)d}:Q\subseteq K_{j,m}}b_{Q}.$
We note that each function $H_{j}^{s}*B^{j-s,m}$ is supported in the cube $2K_{j,m}$. It is easy to see that for each fixed $j\in{\mathbb{Z}}$, the cubes $\{2K_{j,m}\}_{m\in{\mathbb{Z}}^{d}}$ have bounded overlap but they lack the nesting structure. We will further subdivide the cubes $\{K_{j,m},j\in{\mathbb{Z}},m\in{\mathbb{Z}}^{d}\}$ into finitely many baskets based on the shifted dyadic grids defined as following. For $u\in\{0,\frac{1}{3},\frac{2}{3}\}^{d}$, we define the dyadic grids $\mathcal{D}_{u}$ as
$\mathcal{D}_{u}=\{2^{j}([0,1]^{d}+m+u),\;j\in{\mathbb{Z}},\;m\in{\mathbb{Z}}^{d}\}.$
We will require the following fact from [HLP2013, Lemma 2.5].
Lemma 2.1 ([HLP2013]).
[…]
By the above lemma, for each $K_{j,m}$, there exists a $u\in\left\{0,\frac{1}{3},\frac{2}{3}\right\}^{d}$ and $\widetilde{K}_{j,m}\in\mathcal{D}_{u}$ such that $2K_{j,m}\subset\widetilde{K}_{j,m}$ and $|\widetilde{K}_{j,m}|\lesssim|{K}_{j,m}|$. Based on this observation, we have the following decomposition:
$B^{j-s,m}=\sum_{u\in\{0,\frac{1}{3},\frac{2}{3}\}^{d}}B^{j-s,m}_{u},\;\text{where}$
$B^{j-s,m}_{u}=\begin{cases}B^{j-s,m},&\;\text{if}\;\widetilde{K}_{j,m}\in\mathcal{D}_{u}\\ 0,&\;\text{otherwise}.\end{cases}$
Thus, it is enough to estimate
$\left|\left\{V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>{c_{1}3^{-d}2^{-\delta_{1}s}\alpha}\right\}\right|,$
for a fixed $u\in\left\{0,\frac{1}{3},\frac{2}{3}\right\}^{d}$. The advantage of the above expression is that $\operatorname{supp}(H_{j}^{s}*B^{j-s}_{u})\subset\widetilde{K}_{j,m}$, where $\widetilde{K}_{j,m}\in\mathcal{D}_{u}$ and any such two $\widetilde{K}_{j,m}$ and $\widetilde{K}_{j^{\prime},m^{\prime}}$ are either disjoint or one is contained in another. We will now discard the points where there is a large interaction on the scales $j$. This leaves us with points having a controlled overlap in the scales with respect to the parameter $s$. To do so, we initially set
$\mathcal{C}_{0}^{u}$ $=\{(j,m):\exists Q\in\mathcal{Q},|Q|=2^{(j-s)d},\;Q\subseteq K_{j,m}\;\text{and}\;\widetilde{K}_{j,m}\in\mathcal{D}_{u}\},\;\text{and}$
$\mathcal{X}^{u}$ $=\bigcup\limits_{(j,m)\in\mathcal{C}^{u}_{0}}\widetilde{K}_{j,m}.$
It is easy to see that $\operatorname{supp}(\sum_{j\in{\mathbb{Z}}}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u})\subseteq\mathcal{X}^{u}$. We decompose the set $\mathcal{X}^{u}=\mathcal{X}^{u}_{1}+\mathcal{X}^{u}_{2}$, where
$\mathcal{X}^{u}_{1}=\left\{x\in\mathcal{X}^{u}:\sum_{(j,m)\in\mathcal{C}^{u}_{0}}\chi_{\widetilde{K}_{j,m}}(x)>2^{s(d+\epsilon_{0})}\right\},\quad\text{and}\quad\mathcal{X}^{u}_{2}=\mathcal{X}^{u}\setminus\mathcal{X}^{u}_{1},$
and $\epsilon_{0}>0$ is a small fixed constant. Here, we would like to point out that the choice of the sublevel set $\mathcal{X}^{u}_{1}$ based on $\epsilon_{0}>0$ allows us to do away with the recursion argument employed in [KL2020]. We write
$\left|\left\{V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>c_{1}3^{-d}2^{-\delta_{1}s}\alpha\right\}\right|$
$\lesssim$ $|\mathcal{X}^{u}_{1}|+\left|\left\{x\in\mathcal{X}^{u}_{2}:V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>c_{1}3^{-d}2^{-\delta_{1}s}\alpha\right\}\right|.$
We note that for any $(j,m)\in\mathcal{C}^{u}_{0}$, there exists atleast one cube $Q_{j.m}\in\mathcal{Q}$ with $|Q_{j,m}|=2^{(j-s)d}$ and $Q_{j,m}\subset K_{j,m}$. Thus, by Chebyshev’s inequality, we have
$|\mathcal{X}^{u}_{1}|\leq\frac{1}{2^{s(d+\epsilon_{0})}}\sum_{(j,m)\in\mathcal{C}^{u}_{0}}|\widetilde{K}_{j,m}|\lesssim\frac{1}{2^{s\epsilon_{0}}}\sum_{(j,m)\in\mathcal{C}^{u}_{0}}|Q_{j,m}|\lesssim\frac{2^{-s\epsilon_{0}}}{\alpha}\sum\|b_{Q}\|_{1}.$
Let $\mathcal{C}^{u}=\{(j,m)\in\mathcal{C}_{0}^{u}:\widetilde{K}_{j,m}\nsubseteq\mathcal{X}^{u}_{1}\}$. Then, it is easy to see that for $x\in\mathcal{X}^{u}_{2}$, we have
$\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}}H_{j}^{s}*B^{j-s,m}_{u}(x)=\sum_{(j,m)\in\mathcal{C}^{u},\;j\geq k}H_{j}^{s}*B^{j-s,m}_{u}(x).$
The advantage of working with the collection $\mathcal{C}^{u}$ is that for any fixed $(j,m)\in\mathcal{C}^{u}$, there are atmost $2^{s(d+\epsilon_{0})}$ indices $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j,m}\subset\widetilde{K}_{j^{\prime},m^{\prime}}$. Based on this observation, we subdivide the collection $\mathcal{C}^{u}$ as follows. For any $(j,m)\in\mathcal{C}^{u}$, we say $(j,m)\in\mathcal{C}^{u}_{1}$ if there is no other index $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subseteq\widetilde{K}_{j,m}$. Inductively, we define $\mathcal{C}^{u}_{l}$ to be the collection of all indices $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{u}_{i}$ such that there does not exist any index $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{u}_{i}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subset\widetilde{K}_{j,m}$. By the observation before, the process terminates at the step $l=2^{s(d+\epsilon_{0})}$ and we have $\mathcal{C}^{u}=\bigcup\limits_{l=1}^{2^{s(d+\epsilon_{0})}}\mathcal{C}^{u}_{l}$. Thus, we have that
$\left|\left\{x\in\mathcal{X}^{u}_{2}:V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>c_{1}3^{-d}2^{-\delta_{1}s}\alpha\right\}\right|$
$\lesssim\frac{2^{2\delta_{1}s}}{\alpha^{2}}\left\|V_{2}\left(\sum_{(j,m)\in\mathcal{C}^{u}}H_{j}^{s}*B^{j-s,m}_{u}:k\in{\mathbb{Z}}\right)\right\|_{2}^{2}$
$=\frac{2^{2\delta_{1}s}}{\alpha^{2}}\left\|V_{2}\left(\sum_{i=1}^{l}\left(\sum_{(j,m)\in\mathcal{C}^{u}_{i}}H_{j}^{s}*B^{j-s,m}_{u}\right):1\leq l\leq 2^{s(d+\epsilon_{0})}\right)\right\|_{2}^{2}$
$\lesssim\frac{s^{2}2^{2\delta_{1}s}}{\alpha^{2}}\sup_{\varepsilon_{l}\in\{-1,0,1\}}\left\|\sum_{l=1}^{2^{s(d+\epsilon_{0})}}\varepsilon_{l}\left(\sum_{(j,m)\in\mathcal{C}^{u}_{l}}H_{j}^{s}*B^{j-s,m}_{u}\right)\right\|_{2}^{2}$
$=\frac{s^{2}2^{2\delta_{1}s}}{\alpha^{2}}\sup_{\varepsilon_{l}\in\{-1,0,1\}}\left\|\sum_{j\in{\mathbb{Z}}}H_{j}^{s}*\left(\sum_{l=1}^{2^{s(d+\epsilon_{0})}}\sum_{m:(j,m)\in\mathcal{C}^{u}_{l}}\varepsilon_{l}B^{j-s,m}_{u}\right)\right\|_{2}^{2}$
$\lesssim\frac{s^{2}2^{-s(\delta-2\delta_{0}-2\delta_{1})}}{\alpha}\left\|\Omega\right\|_{1}\sum_{Q\in\mathcal{Q}}\|b_{Q}\|_{1},$
where we used Lemma A.2 in the last step. Choosing $\delta_{0},\delta_{1}\leq\frac{\delta}{100}$ and summing in $u$ and $s$ concludes the proof.
Appendix A Microlocal estimates for rough singular integrals
A.1. Estimates of Seeger:
Let $H_{0}$ be a function such that
  • $\operatorname{supp}(H_{0})\subseteq\{2^{-1}\leq|x|\leq 2\}$.
  • For $N\in{\mathbb{N}}$, we have
    $\sup_{0\leq l\leq N}r^{d+l}|\frac{\partial^{l}}{\partial r^{l}}H_{0}(r\theta)|\leq\mathscr{M},$
    uniformly in $r$ and $\theta$.
We set $H_{j}(x)=2^{-jd}H_{0}(2^{-j}x)$. For $s\geq 5$, let $\mathscr{E}^{s}=\{e_{v}^{s}\}$ be a collection of unit vectors such that
  • $|e_{v_{1}}^{s}-e_{v_{2}}^{s}|>2^{-s-10}d^{-1}$ for all $e_{v_{1}}^{s},e_{v_{2}}^{s}\in\mathscr{E}^{s}$.
  • For each $\theta\in{\mathbb{S}}^{d-1}$, there exists $e_{v}^{s}\in\mathscr{E}^{s}$ such that $|\theta-e_{v}^{s}|\leq 2^{-s-1}$.
It is easy to see that there exists collection of disjoint measurable sets $E_{v}^{s}\subset{\mathbb{S}}^{d-1}$ such that $e_{v}^{s}\in E_{v}^{s}$, $\sigma(E_{v}^{s})\leq 2^{-s(d-1)}$ and $\cup_{v}E_{v}^{s}={\mathbb{S}}^{d-1}$. Let $H_{j,v}^{s}(x)=H_{j}(x)\chi_{E_{v}^{s}}(x/|x|)$. For a fixed parameter $\kappa$ with $0<\kappa<1$, we have the following decomposition from[Seeger1996, (2.2)],
$H_{j}=\Gamma_{j}^{s}+(H_{j}-\Gamma_{j}^{s}),$
where $\widehat{\Gamma_{j}^{s}}(\xi)=\sum_{v}\psi(2^{s(1-\kappa)}\langle\xi,e_{v}^{s}\rangle/|\xi|)\widehat{H_{j,v}^{s}}(\xi)$ and $\psi\in C_{c}^{\infty}({\mathbb{R}})$ is supported in [4,4][-4,4] and $\psi(t)=1$ for $t\in[-2,2]$.
We have the following estimates, which is essentially an amalgamation of Lemmas 2.1 and 2.2 in [Seeger1996].
Lemma A.1 ([Seeger1996]).
Let $\alpha>0$ and $\mathcal{Q}$ be a collection of cubes with disjoint interior such that for each $Q\in\mathcal{Q}$ there exists a function $b_{Q}\in L^{1}(Q)$ such that
$\int|b_{Q}(x)|dx\leq\alpha|Q|.$
Let $B^{k}=\sum_{Q\in\mathcal{Q}:|Q|=2^{dk}}b_{Q}$. Then, for $s\geq 5$, we have
$\left\|\sum_{j\in{\mathbb{Z}}}\Gamma_{j}^{s}*B^{j-s}\right\|_{2}^{2}\lesssim\mathscr{M}^{2}2^{-s(1-\kappa)}\alpha\sum_{Q\in\mathcal{Q}}\|b_{Q}\|_{1}.$
Moreover, if $\int b_{Q}=0$ for each $Q\in\mathcal{Q}$. Then, for $N\geq d+1$ and $0\leq\kappa^{\prime}\leq 1$, we have
$\left\|\sum_{j\in{\mathbb{Z}}}(H_{j}-\Gamma_{j}^{s})*B^{j-s}\right\|_{1}\lesssim\mathscr{M}(2^{-s\kappa^{\prime}}+2^{s(d+(\kappa^{\prime}-\kappa)N)})\sum_{Q\in\mathcal{Q}}\|b_{Q}\|_{1}.$
We combine the $L^{1}$ and $L^{2}$ estimates in Lemma A.1 into a single $L^{2}$ estimate in the following manner.
Lemma A.2.
Let $\alpha>0$ and $\mathcal{Q}$ be a collection of cubes with disjoint interior such that for each $Q\in\mathcal{Q}$ there exists a function $b_{Q}\in L^{1}(Q)$ such that
$\int|b_{Q}(x)|dx\leq\alpha|Q|.$
Let $B_{k}=\sum_{Q\in\mathcal{Q}:|Q|=2^{dk}}b_{Q}$. Then, there exists $\delta>0$ such that for $s\geq 5$, we have
$\left\|\sum_{j\in{\mathbb{Z}}}H_{j}*B^{j-s}\right\|_{2}^{2}\lesssim\mathscr{M}^{2}2^{-s\delta}\alpha\sum_{Q\in\mathcal{Q}}\|b_{Q}\|_{1}.$
Proof.
We write
$\sum_{j\in{\mathbb{Z}}}H_{j}*B^{j-s}=\sum_{j\in{\mathbb{Z}}}\Gamma_{j}^{s}*B^{j-s}+\sum_{j\in{\mathbb{Z}}}(H_{j}-\Gamma_{j}^{s})*B^{j-s}=:A+B.$
We observe that
$\int|(A+B)(x)|^{2}dx$
$=$ $\left(\int_{|B(\cdot)|\leq 2^{\rho s}\alpha}+\int_{|A(\cdot)|>2^{\rho s-1}\alpha}+\int_{|(A+B)(\cdot)|>2^{\rho s-1}\alpha}\right)|(A+B)(x)|^{2}dx$
$=:$ $I+II+III.$
Estimate for I: By Lemma A.1 with $\kappa=\frac{1}{2},\;\kappa^{\prime}=\frac{1}{4},\;N=4d+1$, we have
$I$ $\lesssim\int|A(x)|^{2}dx+\int_{|B(\cdot)|\leq 2^{\rho s}\alpha}|B(x)|^{2}dx$
$\lesssim\int|A(x)|^{2}dx+2^{\rho s}\alpha\int_{|B(\cdot)|\leq 2^{\rho s}\alpha}|B(x)|dx$
$\lesssim\mathscr{M}^{2}2^{-\frac{s}{8}}\alpha\sum\|b_{Q}\|_{1},$
where we used $\rho$ such that $\rho<\frac{1}{8}$. To estimate $II$ and $III$, we will need the following claim:
$\left|\left\{x\in{\mathbb{R}}^{d}:|\sum_{j\in{\mathbb{Z}}}H_{j}*B^{j-s}|>\lambda\alpha\right\}\right|\lesssim\frac{\mathscr{M}}{\lambda^{3}\alpha}\sum\|b_{Q}\|_{1},\quad\text{for}\;\lambda>0.$ (A.1)
Assuming the above estimate, we estimate $II$ and $III$.
Estimate for $II$: By Chebyshev’s inequality, Lemma A.1 ($\kappa=\frac{1}{2}$) and (A.1), we have
$II$ $\lesssim\int_{0}^{\infty}\min\left\{\left|\left\{|A(x)|>2^{\rho s-1}\alpha\right\}\right|,\left|\left\{|(A+B)(x)|>\lambda\right\}\right|\right\}\lambda d\lambda$
$\lesssim\int_{0}^{2^{\rho s}\alpha}\left|\left\{|A(x)|>2^{\rho s-1}\alpha\right\}\right|\lambda d\lambda+\int_{2^{\rho s}\alpha}^{\infty}\left|\left\{|(A+B)(x)|>\lambda\right\}\right|\lambda d\lambda$
$\lesssim\|A\|_{2}^{2}+\mathscr{M}\alpha^{2}\sum\|b_{Q}\|_{1}\int_{2^{\rho s}\alpha}^{\infty}\frac{d\lambda}{\lambda^{2}}$
$\lesssim\mathscr{M}^{2}\alpha 2^{-\frac{s}{8}}\sum\|b_{Q}\|_{1}.$
Estimate for $III$: It is not difficult to see that an application of (A.1) implies that
$II$ $\lesssim\int_{0}^{\infty}\left|\left\{|(A+B)(x)|>\max\{\lambda,2^{\rho s-1}\alpha\}\right\}\right|\lambda d\lambda\lesssim\mathscr{M}^{2}\alpha 2^{-\frac{s}{8}}\sum\|b_{Q}\|_{1}.$
We now prove (A.1). We choose a smooth and positive function $\nu_{0}$ supported in $\{2^{-2}\leq|x|\leq 2^{2}\}$ such that $\|\nu_{0}\|_{\infty},\|\nu_{0}\|_{1}\leq 10^{d}\mathscr{M}$ and $H_{0}(x)\leq\nu_{0}(x)$. We will prove (A.1) for $\nu_{j}(x)=2^{-jd}\nu(2^{-j}x)$ instead of $H_{j}$ and (A.1) will be a consequence. We note that the estimate (A.1) follows from the following two estimates
$\left\|\sum_{j\in{\mathbb{Z}}}\nu_{j}*B^{j-s}\right\|_{1}\lesssim\mathscr{M}\sum\|b_{Q}\|_{1},$ (A.2)
$\left\|\sum_{j\in{\mathbb{Z}}}\nu_{j}*B^{j-s}\right\|_{BMO}\lesssim\mathscr{M}\alpha.$ (A.3)
Indeed, by Chebyshev’s inequality and interpolation, we have that
$\left|\left\{x\in{\mathbb{R}}^{d}:|\sum_{j\in{\mathbb{Z}}}H_{j}*B^{j-s}|>\lambda\alpha\right\}\right|$ $\lesssim\frac{1}{(\lambda\alpha)^{3}}\left\|\sum_{j\in{\mathbb{Z}}}\nu_{j}*B^{j-s}\right\|_{3}^{3}$
$\lesssim\frac{\mathscr{M}}{(\lambda\alpha)^{3}}\left[\left(\sum\|b_{Q}\|_{1}\right)^{\frac{1}{3}}\alpha^{\frac{2}{3}}\right]^{3}$
$\lesssim\frac{\mathscr{M}}{\lambda^{3}\alpha}\sum\|b_{Q}\|_{1}.$
The estimate (A.2) is easy to verify.
We will provide the proof for inequality (A.3). First, we observe that by the disjointness of the collection $\mathcal{Q}$, for any $F\subset{\mathbb{R}}^{d}$, we have
$\sum_{j\in{\mathbb{Z}}}\|B^{j-s}\chi_{F}\|_{1}\sum_{Q\in\mathcal{Q}:Q\cap F\neq\emptyset}\alpha|Q|\leq\alpha|F|.$ (A.4)
Now, we fix a cube $R\subset{\mathbb{R}}^{d}$, we write
$\frac{1}{|R|}\int_{R}\left|\sum_{j\in{\mathbb{Z}}}\nu_{j}*|B^{j-s}|(x)-\frac{1}{|R|}\int_{R}\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(y)dy\right|dx$
$\leq$ $\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(x)-\frac{1}{|R|}\int_{R}\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(y)dy\right|dx$
$+\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}<\ell(R)}\nu_{j}*|B^{j-s}|(x)\right|dx$
$=:$ $S_{1}+S_{2}.$
By mean value theorem and local support of the functions, we have
$S_{1}$ $\lesssim\frac{1}{|R|}\int_{R}\sum_{2^{j}\geq\ell(R)}2^{-j}\ell(R)\|(\nabla\nu)_{j}\|_{\infty}\|B^{j-s}\chi_{B(c_{R},2^{j+3})}|\|_{1}dx$
$\lesssim\sum_{2^{j}\geq\ell(R)}2^{-j}\ell(R)2^{-jd}\mathscr{M}\alpha|B(c_{R},2^{j+3})|$
$\lesssim\mathscr{M}\alpha,$
where $C_{R}$ is the center of the cube $R$ and we used (A.4) in the second step.
By the local support of the functions and (A.4), we have
$S_{2}$ $\lesssim\frac{1}{|R|}\sum_{2^{j}<\ell(R)}\|\nu_{j}\|_{1}\|B^{j-s}\chi_{6R}\|_{1}$
$\lesssim\frac{\mathscr{M}}{|R|}\sum_{j\in{\mathbb{Z}}}\|B^{j-s}\chi_{6R}\|_{1}$
$\lesssim\frac{\mathscr{M}}{|R|}\alpha|R|=\mathscr{M}\alpha.$
Thus, (A.3) follows and this concludes the proof. ∎
Acknowledgement
Ankit Bhojak would like to thank Michael Lacey and Parasar Mohanty for various helpful conversations. Ankit Bhojak is supported by the Science and Engineering Research Board, Department of Science and Technology, Govt. of India, under the scheme National Post-Doctoral Fellowship, file no. PDF/2023/000708.